Without doing the Gram-Schmidt process again, write a QR decomposition for A = [2 7 14 2 4 - 7 1 5 41 5 4 A= 2 4 2 4 -7 2 7 14
Answer #1

The given matrix is: A= 2.00 7.00 14.00 2.00 4.00 -7.00 1.00 5.00 4.00 k = 0 ao 2.00 2.00 1.00 -0.67 -0.67 -0.33 Но -0.67 0.7-0.67 0.33 -0.67 -0.67 -0.67 0.33 Q1 -0.33 0.67 0.67 -3.00 -9.00 -6.00 A1 0.00 3.00 12.00 0.00 0.00 -9.00 a = 1.25 B = 0.36 kResult of the QR decomposition Q -0.67 0.33 0.67 -0.67 -0.67 -0.33 -0.33 0.67 -0.67 -3.00 -9.00 -6.00 0.00 3.00 12.00 R= 0.00 this is QR decomposition for A' by Householder method. Next for A.

The given matrix is: A 1.00 5.00 4.00 2.00 4.00 -7.00 2.00 7.00 14.00 k = 0 1.00 ao = 2.00 2.00 Но -0.33 -0.67 -0.67 -0.67 0.-0.33 0.67 -0.67 -0.67 -0.67 -0.33 Q1 = -0.67 0.33 0.67 -3.00 -9.00 -6.00 0.00 3.00 12.00 A1 - 0.00 0.00 9.00 a = 1.00 B = 0.Result of the QR decomposition -0.33 0.67 0.67 Q = -0.67 -0.67 0.33 -0.67 0.33 -0.67 -3.00 -9.00 -6.00 R ( 0.00 3.00 12.00 -9 this is of A.